OpenStudy (faman39):

What is the experimental probability of rolling a 5 on a standard die if you had 50 trials and 15 successes? 3% 15% 30% 35%

OpenStudy (anonymous):

P = 15/ 50 = 3%

OpenStudy (anonymous):

no. of times 5 is obtained/total no. of attempts. in this case, 15/50=3/10